A Simple pendulum is suspended from the ceiling of a lift. When the lift Is at rest, )Its time period ls T.Wlth what acceleration should lift be accelerated upward in order to ` reduce its time period to T/2 ? (1) g (2) 2g(3)3g(4)-3g

Grade:11

1 Answers

Rahul Verma
7 years ago
From the equation of time period of simple pendulum,
T=2\pi \sqrt{\frac{l}{g}}      ______________(1)
we get acc to quesstion
\frac{T}{2}=2\pi \sqrt{\frac{l}{g+a}}          ________________(2)
dividing 1 and 2 we get acceleration equal to g